OpenStudy (anonymous):

log 2^(2+1)=4 -log2^(x+7)

OpenStudy (dumbcow):

is that \[\log 2^{2x+1} = 4-\log 2^{x+7}\]

OpenStudy (dumbcow):

??

OpenStudy (anonymous):

yes

OpenStudy (dumbcow):

anytime you have an exponent inside a log, you can bring it out front \[(2x+1)\log 2 = 4 - (x+7)\log 2\] distribute and combine like terms \[(3\log 2)x = 4-8\log 2\] divide \[x = \frac{4-8\log 2}{3\log 2}\]